im a right triangle, find the length of the side not given.
b=2, c-square root of 29

Answer by ewatrrr(24785)


Hi,
Let a represent the length of the side not given
Applying the Pythagoream Theorem
a^2 + 2^2 = 29
a^2 = 25
a = 5
CHECKING our Answer
25 + 4 = 29